Circular No. 5616

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1E 1740.7–2942 B. Cordier, Service d’Astrophysique, Saclay; M. C. Schmitz-Fraysse,

Centre d’Etude Spatiale des Rayonnements, Toulouse; and R. Sunyaev and M. Gilfanov, Space Research Institute, Moscow, report for the Granat SIGMA team: “The SIGMA telescope has observed the region of the Galactic Center on Sept. 14-15 for a total exposure live-time of 46 660 s. The soft γ-ray source 1E 1740.7–2942 has been detected at a flux level of 97 ± 12 mCrab in the band 40-150 keV. The source flux appears to be increasing with respect to the 1992 April SIGMA observations, though it has not yet attained the 1990 March/April level. Conversely, the source GRS 1758–258 remains in its low state, since it has not been detected with a 2σ upper limit (40-150 keV) of ∼ 25 mCrab.”

NOVA IN NGC 1313? J. M. Chapman, R. P. Norris, and J. E. Reynolds, Australia Telescope

National Facility, CSIRO; and P. te Lintel Hekkert, Mount Stromlo and Siding Spring Observatories, report that this nova/supernova candidate (cf. IAUC 5615) was unresolved with the Parkes-Tidbinbilla Interferometer at 1.6 GHz from an observation taken in 1992 May, giving an upper limit to its diameter of 0′′.1.

COMET HELIN-LAWRENCE (1992q)

Orbital elements from MPC 20775:

T = 1993 Mar. 11.131 TT  ω = 266.^{o}293
Ω = 194.882 }2000.0
q = 2.09993 AU  i = 107.139
1992TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 25 1^{h}36.^{m}77 −27^{o}57.^{′}4 1.974 2.849 144.^{o}2 11.^{o}9 14.5
30 1 28.38 −31 06.5 1.947 2.812 143.0 12.4 14.4
Oct. 5 1 18.96 −34 10.5 1.933 2.776 140.2 13.3 14.4
10 1 08.62 −37 05.3 1.931 2.740 136.2 14.6 14.3
15 0 57.50 −39 47.1 1.940 2.705 131.3 16.1 14.3
20 0 45.82 −42 13.1 1.961 2.670 125.9 17.6 14.2
25 0 33.84 −44 21.5 1.991 2.636 120.3 19.0 14.2

1992 September 23 Daniel W. E. Green